How large a sample n would you need to estimate p with margin of error 0.04 with 90% confidence? Assume that you don’t know anything about the value of p.

a. n = 601

b. n = 423

c. n = 1068

d. n = 608

e. n = 2305
Answer #1

Solution :

Given that,

= 1 - = 0.5

margin of error = E = 0.04

Z/2 = 1.645

sample size = n = (Z / 2 / E )2 * * (1 - )

= (1.645 / 0.04)2 * 0.5 * 0.5

= 423

sample size = n = 423

option b. is correct